OpenStudy (anonymous):

4m^3/3m^2n times 9n^2/30m^2n^2

OpenStudy (pfenn1):

\[\frac{4m^3}{3m^2n} \times \frac{9n^2}{30m^2n^2}\]Can you simplify any of the terms?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

im lazy

OpenStudy (pfenn1):

You can do it. Just try. Or multiply the terms out and see what that gets you.